OpenStudy (anonymous):

Find, without using a calculator, all the possible values of y between 0 and 2pie for which tan(2y + pie/3 ) = 1

OpenStudy (rsadhvika):

period of tan is pi

OpenStudy (anonymous):

\[\tan(2y+\frac{\pi}{3})=1====>2y+\frac{\pi}{3}=\pi/4 +k \pi\] with k in {.....-1,0,1,2.....} find k wich satisfies y in 0 2pi

OpenStudy (anonymous):

Erm just to verify do you get y = - pie/ 24 , 11pie/ 24 , 23pie/24 , 35pie / 24

OpenStudy (anonymous):

the first one is not incude ! -pi/24<0 :) !

OpenStudy (anonymous):

I beg your pardon?

OpenStudy (anonymous):

we are looking for y beetween 0 and 2pie so ! not all values of k are good ! you have to choose ! which ones satifies y beetween 0 and 2 pi ! Did you understand ?!

OpenStudy (anonymous):

yes i did the which quadrant method and tan y is in the first quadrant thus, basic angle pie/ 4.

OpenStudy (anonymous):

tan y = 1*

OpenStudy (anonymous):

so the answers are ...?!

OpenStudy (anonymous):

i already typed my answers duhhhh above.

OpenStudy (anonymous):

what answer :) ?!

OpenStudy (anonymous):

y = - pie/ 24 , 11pie/ 24 , 23pie/24 , 35pie / 24

OpenStudy (anonymous):

A=2y and B=pie/3 tan (A + B) = (tan A + tan B)/(1 - (tan A)(tan B)) now put the values of A and B

OpenStudy (anonymous):

i hope its easy for you.

OpenStudy (anonymous):

y=-pi/24 don't satisfy the condition y>=0 see ?@Sgstudent

OpenStudy (anonymous):

and you missed one ! @Sgstudent

OpenStudy (anonymous):

are you there ? @Sgstudent

OpenStudy (anonymous):

i am trying muhammad method.

OpenStudy (anonymous):

to be honest this is my first time attempting tan (A+B) i have attempted sin and cos (A+B) Oh welll

OpenStudy (anonymous):

how to do his method lol

OpenStudy (anonymous):

as in i need the workings.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

:D :D :D ! what you didn't understand in ""my method" ?! and also muhammad's methods suppose that 2y==/==pi/2 and tan(2y).tan(pi/3) ==/==1 so you're solving an equation ! just it's inappriopriate to do hypothisis for the unkown :) ! I can show a full answer if you want to ! (there are two )

OpenStudy (anonymous):

anything :)

OpenStudy (anonymous):

i think A+B is easier

OpenStudy (anonymous):

i love solving eqn :) :DD

OpenStudy (anonymous):

I don't think so :) ! A+B is more dangerous !

OpenStudy (anonymous):

my brain is different oh yea

OpenStudy (anonymous):

did you understand the first part ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i dont know how to get tan 2 y values

OpenStudy (anonymous):

let cange the problem ! Can you solve the equation \[\tan(\alpha)=1\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x= tan inverse

OpenStudy (anonymous):

that's strange ! you said that you love solving equations ! :)

OpenStudy (anonymous):

x = tan inverse wut/

OpenStudy (anonymous):

kk I'll make you understand :) wait a minute

OpenStudy (anonymous):

i just solved a tedious quadratic equation haha :P

OpenStudy (anonymous):

look at the attachement ! to understand tan(x)=1 first !

OpenStudy (anonymous):

quadratic equations are really boring :) !

OpenStudy (anonymous):

done understanding

OpenStudy (anonymous):

but trig equations are fascinating ! (sometimes not always !) did you understand ! ?

OpenStudy (anonymous):

i understandddddddddddddddddddd

OpenStudy (anonymous):

are you angry ?! or what ! now do you agree with tan(smthing)=1<=====>smthing=pi/4+ k pi

OpenStudy (anonymous):

i dont have anger management prob just i understand already :P

OpenStudy (anonymous):

we continue ?

OpenStudy (anonymous):

yes sure "_"

OpenStudy (anonymous):

what does this face mean ?! we said that tan (smthing)=1<=====>smthing=pi/4+k pi (wich you understood certainly) ! \[\tan(2y+\frac{\pi}{3})=1<=====>2y+1= \frac{\pi}{4}+k \pi \] with k is an integer (this part is cleaar )

OpenStudy (anonymous):

this face means that i am happy to learn more

OpenStudy (anonymous):

:D :D you're making me happy ! then !

OpenStudy (anonymous):

is it clear shall we continue ?

OpenStudy (anonymous):

yes yes keep it going

OpenStudy (anonymous):

\[2y=-\frac{\pi}{3}+\frac{\pi}{4}+k \pi =- \frac{\pi}{12}+ k \pi =\frac{12k \pi-\pi}{12}\] then \[y=\frac{12k \pi -\pi }{24}\] is this part cleaar .? ( now I need all your mind )

OpenStudy (anonymous):

FOCUS*

OpenStudy (anonymous):

now we will discuss about this kk ?§

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

now they want us to find y beetween 0 an 2pi now we have to choose k ! if we choose k=-1 (for example) is " y valid ?! is y beetween 0 an 2pi !

OpenStudy (anonymous):

please don't leaave me alone ! :):)

OpenStudy (anonymous):

I am here :(

OpenStudy (anonymous):

you are not alone , i am here with you , though you far away i am here to stay lol quoted by MJ

OpenStudy (anonymous):

ohh ! I was checking your profile :) ! now :) Answer the question !

OpenStudy (anonymous):

What was the question?

OpenStudy (anonymous):

now they want us to find y beetween 0 an 2pi now we have to choose k ! if we choose k=-1 (for example) is " y valid ?! is y beetween 0 an 2pi !

OpenStudy (anonymous):

it is unvalid because it cant be <0 theeeeeeeeee

OpenStudy (anonymous):

are you angry again ? now k=0

OpenStudy (anonymous):

i am not angry i am very happy LOLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLL

OpenStudy (anonymous):

k = 0 also cannot :(

OpenStudy (anonymous):

k=1 ; k=3; k=4 ;k=5,k=6( now ! I bet you're not happy )

OpenStudy (anonymous):

k=2 :D don't forget

OpenStudy (anonymous):

nonsense LOL

OpenStudy (anonymous):

the second method a lot easier \[ 0 \le y \le 2 \pi \Rightarrow 0 \le \frac{12k \pi -\pi}{24} \le 2 \pi\] which Implies \[0 \le \frac{12k-1}{24} \le 2 \] then \[1 \le 12 k \le 49\] \[\frac{1}{12}\le k \le \frac{49}{12}\] because k is an integer k =0 or 1 or 2 or 3 or 4 then you'll have 5 solutions ! I hope it's clear now !

OpenStudy (anonymous):

are you there Hapyy man or woman ?!